• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name LeRose

    Log entry time 16:07:58 on October24,2002

    Entry number 88951

    keyword=Shift summary

    Day shift 10/24/02
    John LeRose (SL)
    Kent Paschke (TO)
    Salvatore Frullani

    shift starts with Q1 left down. Requires access to reset (software reset doesn't work!)
    8:34 Beam back started run 2727 (Q3D_d50 production run)
    9:45 start run 2728
    10:10 Run ended. MCC taking beam for 10 minutes
    10:30 MCC starts trying to give us beam.
    11:03 beam is back start run 2729
    11:11 MCC says Delivering beam is interfering with the machine development program so we'll have no beam for a while. Run 2729 ended.
    12:20 MCC says they are ready to start giving us CW beam. Arun Saha informed that he can start his arc energy measurement
    12:30 NOW they say they're no quite ready, they're still doing some tests. They'll call us when they really are ready.
    13:00 MCC now says they have stable beam and are ready to start setup for ARC energy measurement (paged Arun)
    13:08 finally got Arun in his office. He's calling MCC
    13:09 MCC (Sandy) called again. They're still ready. I suggested they call Arun.
    14:07 Arun is done (see halog 88948). MCC backing out, i.e. returning to acromatic mode.
    14:30 Beam returns (but not really)
    15:08 MCC says they're having trouble with their orbits
    15:40 MCC says the machine is back in order but now they have to make some minor adjustments to our arc.
    15:45 They start trying to give us CW
    16:00 They say they have trouble getting from 100 to 110 microamps we decide to go with 100
    executive summary: 8 hours clock time, 2.45 hours ABU, 12 good events (1 accidental?)